# Grasping the Essentials: Tailoring Large Language Models for Zero-Shot Relation Extraction

### v1

@@ Abstract
Relation extraction (RE), a crucial task in NLP, aims to identify semantic relationships between entities mentioned in texts. Despite significant advancements in this field, existing models typically rely on extensive annotated data for training, which can be both costly and time-consuming to acquire. Moreover, these models often struggle to adapt to new or unseen relationships. In contrast, few-shot learning settings, which aim to reduce annotation requirements, may offer incomplete and biased supervision for understanding target relation semantics, leading to degraded and unstable performance. To provide the model with accurate and explicit descriptions of the relations types and meanwhile minimize the annotation requirements, we study the definition only zero-shot RE setting where only relation definitions expressed in natural language are used to train a RE model. Motivated by the strong synthetic data generation power of LLMs, we propose a framework REPaL which consists of three stages: (1) We utilize LLMs to generate initial seed instances based on relation definitions and an unlabeled corpora. (2) We fine-tune a bidirectional Small Language Model (SLM) using these initial seeds to learn the relations for the target domain. (3) We enhance pattern coverage and mitigate bias resulting from the limited number of initial seeds by incorporating feedback acquired from SLM’s predictions on unlabeled corpora. To accomplish this, we leverage the multi-turn conversation ability of LLMs to generate new instances in follow-up dialogues. Experiments on two datasets show REPaL achieves better zero-shot performance with large margins over baseline methods.
@@
Sizhe Zhou1, Yu Meng2, Bowen Jin1, Jiawei Han1 1 University of Illinois Urbana-Champaign  2 University of Virginia {sizhz, bowenj4, hanj}@illinois.edu, yumeng5@virginia.edu
Figure 1: Different types of initial seeds for low-resource RE approaches. The relation is based on P276 of FewRel Dataset (Han et al., 2018b). For instance-level seeds, we can see two instance examples fail to cover structure type entity mentions of <ENT0>.
@@ 1 Introduction
Relation extraction is an important Information Extraction (IE) task that aims to identify and classify semantic relationships between entities in natural language texts. It has wide applications in various downstream tasks including knowledge graph construction (Lin et al., 2015; Yu et al., 2020), question answering (QA) (Wang et al., 2012, 2016) and event mining (Jiao et al., 2022; Li et al., 2023c).
Existing works propose to train RE models with massive human-labeled data (Han et al., 2018a, 2020; Yamada et al., 2020). However, in real-world scenarios, it is generally challenging and costly to obtain large-scale and high-quality datasets. To tackle such data scarcity issues, recent works explore few- or zero-shot relation extraction settings. Given that pre-trained language models (PLMs) (Radford et al., 2018; Devlin et al., 2018) store factual relation knowledge (Petroni et al., 2019), the first category of methods proposes to directly prompt them with human-crafted patterns (Zhou et al., 2023a; Sainz et al., 2021; Chia et al., 2022; Han et al., 2022; Chen et al., 2022a). Another category of methods fine-tunes PLMs to capture the relation semantics from different types of texts (Gao et al., 2020; Zhao et al., 2023).
More recently, the remarkable reasoning and generation power of LLMs make it possible to attain better RE performance with fewer human annotations: Some methods exploit the in-context learning paradigm or assistance from other language models to enhance semantic understanding of target relations (Wan et al., 2023; Ma et al., 2023). Other works conduct relational data synthesis to mitigate noise or hallucination issues (Chia et al., 2022; Li et al., 2023b; Sun et al., 2024).
Despite the unprecedented advance achieved by PLM/LLM-based few-shot or zero-shot relation extraction works, they suffer from two issues. The first issue is the underutilization of relation definitions, which offer explicit and accurate relation semantics and are less biased than a limited number of demonstrations. Specifically, many few-shot or zero-shot RE approaches start with instance-level seeds such as relation instances and relation triples (see Fig. 1). Though simple and versatile, such seeds occasionally result in partial coverage of target relation semantics, thus the resulting RE models may fail to grasp the complete relational semantics. Unlike other tasks such as text classification where label semantics are relatively simple (Meng et al., 2020), relation semantics are generally directional and multifaceted which involve entity-entity interactions and entity-related requirements (see Fig. 1). Such relation semantic complexity requires detailed elaborations and is considered in the annotation process of most RE datasets  [^1] For example, SemEval-2010 Task 8 dataset guideline and TAC KBP guidelines but frequently gets neglected in RE model developments.
Another issue is that the potential of LLMs for zero-shot relation extraction is still far from being fully exploited. First, while LLMs are primarily developed and extensively pre-trained for domain adaptation (Radford et al., 2019), current LLM-based zero-shot RE approaches often struggle to achieve satisfactory performance. This suggests that the zero-shot learning capabilities of LLMs may not have been fully utilized. Second, most LLMs are designed to perform multi-turn conversations and excel in seeking feedback from the dialogue history. Such a feature has shown great potential in a vast range of other NLP tasks including knowledge-intensive or complex reasoning question-answering tasks (Trivedi et al., 2022; Zhou et al., 2023b, c). Nevertheless, zero-shot and few-shot LLM-based RE approaches mainly perform the task with single-turn usages of LLMs.
@@
In this work, we focus on definition-based low-resource relation extraction. We study a new zero-shot RE setting where only relation definitions, instead of relation instances, are provided. In addition to the fact that relation definitions serve as less biased initial seeds, such a task setup is realistic as: (1) many downstream applications such as QA tasks already have explicit definitions of interested relations and obtaining such supervision is generally straightforward; and (2) such a setting highlights the importance for RE systems to continuously adapt to new relation types based on corresponding definitions without maintaining a large amount of seen instances and re-training models.
To better leverage LLMs for the proposed zero-shot RE setting, we propose a novel zero-shot RE framework, REPaL. Firstly, REPaL prompts LLMs to generate positive instances conditioned on the definition and samples negative instances from the unlabeled corpus, thus constructing an initial training set. We then train a bidirectional SLM on these instances for training and inference efficiency.
@@
Secondly, to improve pattern coverage of instance generation and rectify the bias within the generated instances and the trained SLM, REPaL acquires instance-level feedback from the trained SLM’s inference results on the unlabeled corpus. Such feedback is utilized to: (1) leverage LLMs’ multi-turn conversational ability to recognize the pattern coverage and existing LLM-SLM’s bias, and then generate instances with new positive patterns and rectified positive patterns, and (2) leverage LLMs’ reasoning ability to diagnose the SLM’s bias and further generate targeted negative instances to rectify such bias. The whole framework performs iterative refinements in which more and better relation instances are accumulated to improve the task-specialized RE model. To foster reproducibility, our data and codes will be organized and available soon.
Our contributions are as follows:
  • We propose the definition only zero-shot RE setting along with modified evaluation datasets to address the deficiency of traditional low-resource RE approaches.
  • We propose a novel zero-shot RE framework that leverages LLMs and unlabeled corpus to generate relation instances and self-improves the semantic coverage while rectifying bias by automatically acquiring and reasoning over feedback.
@@
  • Extensive quantitative and qualitative experiments demonstrate the effectiveness as well as the potential of our proposed task setup and framework.
2 Problem Formulation
@@
Definition 2.1. Definition Only Zero-Shot Relation Extraction. Specifically, we assume that, for any target relation $r(E_{0},E_{1})\in\mathcal{R}(E_{0},E_{1})$  [^2] This work focuses on sentence-level binary relation extraction setting, so two entity mentions are considered for each instance., only one associated relation definition $d(E_{0},E_{1})$ is given. Here $\mathcal{R}(E_{0},E_{1})$ denotes the whole binary relation space and $d(E_{0},E_{1})$ can be a single sentence or a document specifying the target relation $r(E_{0},E_{1})$. $E_{0}$ and $E_{1}$ are two entity placeholders.
@@
The goal of Definition Only Zero-Shot Relation Extraction task is to extract all relation instances that belong to target relation $r(E_{0},E_{1})$ from a given set of relation instances $\{(s^{i},e_{0}^{i},e_{1}^{i})\}_{i=1}^{n}$. Here $s^{i}$ denotes the $i$-th context text, while $e_{0}^{i}$ and $e_{1}^{i}$ denote two entity mentions inside $s^{i}$. Compared with the traditional zero-shot RE task setting, our task assumes no seen data but only target relation definitions. Our task also assumes the unknown negative relation space while mainstream zero-shot RE assumes known information of all test relations. This makes our task more challenging, but aligns better with real-world scenarios.
@@ 3 Methodology
Our proposed REPaL is an iterative framework which consists of three major components: (1) Definition-based seed construction with LLM and the unlabeled corpus (Sec. 3.1); (2) Pattern learning with RE-specialized SLM (Sec. 3.2); and (3) Feedback-driven instance improvement and bias rectification (Sec. 3.3). The overview of REPaL can be found in Fig. 2.
@@ 3.1 Definition-Based Seed Construction with LLM and Unlabeled Corpus
LLMs, such as the GPT family, are pre-trained for the domain adaptation ability (Radford et al., 2019). They have demonstrated to contain factual relation knowledge (Petroni et al., 2019) and are widely used as data generator for various tasks such as text classification (Meng et al., 2022) and dialogue systems (Abdullin et al., 2024). Follow-up evaluation studies have shown that LLMs are relatively skilled at constrained content generation, storytelling, and rationale generation (Sun et al., 2023; Keskar et al., 2019). Following such observations as well as the goal to tackle the data scarcity issue of zero-shot RE, REPaL first leverages LLMs to generate $K_{p_{0}}$ positive seeds, $\{(s_{p}^{i},e_{0,p}^{i},e_{1,p}^{i})\}_{i=1}^{K_{p_{0}}}$, conditioned on the target relation definition $r(E_{0},E_{1})$. We carefully design three prompts for this step (shown in Appendix A) to encourage the LLM to generate comprehensive patterns from three perspectives in terms of complexity: brief, medium, and implicit. Based on exploratory experiments, they yield different types of relation patterns in accordance with our design expectations. Details of experiments and quantitatively how choice of different prompts affect the results are in Sec. 5.3.
In addition to positive seed generation, constructing negative seeds is also vital for better extraction for the target relation $r(E_{0},E_{1})$, as our task setting assumes no prior knowledge on the negative relation space in order to mimic the real-life deployments (Li et al., 2022). Our negative seed construction is based on the hypothesis that, with a normal large-scale unlabeled corpus, the proportion of target relation instances is relatively small. Thus REPaL conducts random sampling over the unlabeled corpus, which is expected to yield an acceptable set of negative instances denoted as $\{(s_{n}^{i},e_{0,n}^{i},e_{1,n}^{i})\}_{i=1}^{K_{n_{0}}}$. In case of extreme cases, REPaL further designs counter-measures in Sec. 3.3.
@@ 3.2 Pattern Learning with RE-Specialized SLM
Given the fact that LLMs are computationally expensive to be fine-tuned, while pure in-context learning (ICL) is hindered by LLM’s sequence length limit, LLM’s capacity to comprehend demonstrations, and LLM’s temporal and monetary expense (Ma et al., 2023), we leverage SLMs as tunable and task-specialized extractors. Specifically, we adapt a natural language inference (NLI) model (Obamuyide and Vlachos, 2018; Liu et al., 2019; Sainz et al., 2021) to be relation-specific binary classifiers for simplicity and leave other architectures for future research.
@@
where $C,N,E$ denote logits $\mathbf{z}$’s indices for NLI label CONTRADICTION, NEUTRAL, ENTAILMENT respectively. And the binary classification loss (Shannon, 1948) for relation $r(E_{0},E_{1})$ is defined as:
@@ 3.3 Feedback-Driven Instance Improvement and Bias Rectification
After obtaining the trained SLM relation extractor, REPaL further introduces a follow-up feedback-driven iterative refinement approach with intuitions listed as follows: (1) the initially generated positive seeds might only have a partial relation pattern coverage. Instead of one-time generation, iterative follow-up generations by incorporating feedback from examining previously generated results are better for pattern coverage. (2) bias could exist after the previous two steps (Sec. 3.1 and 3.2). One potential source of bias is the generated positive seeds or the randomly sampled negative seeds. Another potential source of bias is the training process of SLMs over limited training samples. Identifying and rectifying bias requires a feedback-based follow-up refinement approach.
Specifically, we first obtain the inference results on the unlabeled corpus leveraging trained SLM RE model. The feedback is constructed by sampling instances within certain prediction confidence range. Feedback is constructed for follow-up positive instance generation and for follow-up negative instance generation respectively. The only difference between the above two feedback constructions lies in two distinct prediction confidence ranges.
Secondly, for follow-up positive instance generation, we fully leverage the multi-turn conversational feature of LLMs where the positive instance generation history is maintained and a new turn of conversation asking for positive instances with different relation patterns is appended. The feedback is integrated into the new turn of conversation so that LLMs can identify the correct and incorrect patterns learned by SLMs.
For follow-up negative instance generation, we divide it into two sub-steps: the first sub-step is feedback-driven negative relation definition generation and the subsequent sub-step is negative instance generation based on the negative relation definitions. LLM is queried to examine the constructed feedback and identify the incorrect patterns learned by SLM so that newly generated negative relations can address such patterns. We also query LLM to generate near-miss negative relation definitions to better distinguish hard cases. After acquiring the negative relation definitions, the second sub-step is the same as the initial positive seed generation process in Sec. 3.1. The prompts used in this section are placed in Appendix B. After obtaining all the follow-up positive and negative instances, we iteratively train the SLM RE model.

### v2

@@ Abstract
Relation extraction (RE) aims to identify semantic relationships between entities within text. Despite considerable advancements, existing models predominantly require extensive annotated training data, which is both costly and labor-intensive to collect. Moreover, these models often struggle to adapt to new or unseen relations. Few-shot learning, aiming to lessen annotation demands, typically provides incomplete and biased supervision for target relations, leading to degraded and unstable performance. To accurately and explicitly describe relation semantics while minimizing annotation demands, we explore the definition only zero-shot RE setting where only relation definitions expressed in natural language are used to train a RE model. We introduce REPaL, comprising three stages: (1) We leverage large language models (LLMs) to generate initial seed instances from relation definitions and an unlabeled corpus. (2) We fine-tune a bidirectional Small Language Model (SLM) with initial seeds to learn relations for the target domain. (3) We expand pattern coverage and mitigate bias from initial seeds by integrating feedback from the SLM’s predictions on the unlabeled corpus and the synthesis history. To accomplish this, we leverage the multi-turn conversation ability of LLMs to generate new instances in follow-up dialogues, informed by both the feedback and synthesis history. Studies reveal that definition-oriented seed synthesis enhances pattern coverage whereas indiscriminately increasing seed quantity leads to performance saturation. Experiments on two datasets show REPaL significantly improved cost-effective zero-shot performance by large margins.
@@
Sizhe Zhou1, Yu Meng2, Bowen Jin1, Jiawei Han1 1 University of Illinois Urbana-Champaign  2 University of Virginia {sizhez, bowenj4, hanj}@illinois.edu, yumeng5@virginia.edu
Figure 1: Different types of initial seeds for low-resource RE approaches for example relation P276. It shows using only two instances as seeds fail to cover structure type head entities.
@@ 1 Introduction
Relation extraction is a pivotal task in Information Extraction (IE) that involves identifying and classifying semantic relationships between entities within texts. It has wide applications in various downstream tasks including knowledge graph construction (Lin et al., 2015; Yu et al., 2020), question answering (QA) (Wang et al., 2012, 2016) and event mining (Jiao et al., 2022; Li et al., 2023c). Traditional RE works train models with human-labeled data (Han et al., 2018a, 2020; Yamada et al., 2020). However, acquiring large-scale, high-quality datasets is challenging and costly in reality. To address this data scarcity, few- and zero-shot RE works try to leverage knowledge from LM pre-training or auxiliary sources (Petroni et al., 2019; Chia et al., 2022; Han et al., 2022; Chen et al., 2022a; Zhao et al., 2023; Zhou et al., 2023a; Wan et al., 2023; Li et al., 2023b; Sun et al., 2024).
Despite these advancements, two issues persist in low-resource RE. The first issue is the underutilization of relation definitions. Relation semantics are generally directional and multifaceted which involve entity-entity interactions and entity-related requirements (see Fig. 1). Thus, target relation semantics typically can only be partially reflected by most low-resource supervision, such as seed instances, triples, or label names. Such relation semantic complexity requires detailed elaborations described by relation definitions. Another issue is the underutilization of LLMs for zero-shot RE. Most LLMs are designed to perform multi-turn conversations and excel in seeking feedback from the dialogue history. Such a feature has shown great potential in knowledge-intensive or complex reasoning question-answering tasks (Trivedi et al., 2022; Zhou et al., 2023b, c). Nevertheless, LLM-based low-resource RE works typically rely on single-turn usages.
@@
To address the first issue, this work introduces a new zero-shot RE setting where only relation definitions, instead of seen instances, are provided. In addition to the fact that relation definitions serve as more precise and less biased initial seeds, such a task setup is realistic as: (1) downstream applications such as QA tasks already have explicit definitions of interested relations and obtaining such supervision is generally straightforward; and (2) such a setting highlights the importance for RE systems to continuously adapt to new relation types based on corresponding definitions without maintaining a large amount of seen instances and re-training models.
@@
To address the second issue, we propose a novel zero-shot RE framework, REPaL. REPaL initiates by prompting LLMs to generate positive instances based on predefined definitions and samples negative instances from an unlabeled corpus, forming an initial training set. This set is then used to train an SLM for inference efficiency and performance. Secondly, REPaL acquires and incorporates feedback to address coverage and bias issues from instance generation and SLM training. For robustness, the feedback consists of two independent components: the synthesis dialogues and sampled SLM’s inference results on a large unlabeled corpus. The feedback is utilized to: (1) leverage LLMs’ multi-turn conversational ability to recognize the pattern coverage bias, synthesis error, and then generate instances with new or rectified positive patterns, and (2) leverage LLMs’ reasoning ability to diagnose the SLM’s bias and further generate targeted or near-miss negative instances to rectify such bias by explicitly deriving negative definitions. The whole framework performs iterative refinements in which more and better-quality relation instances are accumulated to improve the task-specialized RE model.
Our data and codes are available here [^1] https://github.com/KevinSRR/REPaL and our contributions are as follows:
  • We demonstrate the partial coverage issue of few-shot RE’s initial seeds. Our studied definition-oriented RE setting can seamlessly leverage few-shot supervision for better pattern coverage and better performance by definition derivation and instance augmentation.
  • We propose a novel zero-shot RE framework, REPaL, that only requires relation definitions and an unlabeled corpus. REPaL iteratively synthesizes both positive and negative instances to enhance pattern coverage and addresses biases by automatically mining and reflecting on feedback from multiple sources, leveraging the multi-turn conversation capability of LLMs.
@@
  • Extensive quantitative and qualitative experiments demonstrate the effectiveness and the potential of our task setup and framework.
2 Background
2.1 Definition Only Zero-Shot Relation Extraction Task
@@
Task Definition 2.1. Definition Only Zero-Shot Relation Extraction. We assume that, for any target relation $r(E_{0},E_{1})\in\mathcal{R}(E_{0},E_{1})$ [^2] This work addresses sentence-level binary relation extraction, where each instance involves evaluating the relationship between two specific entity mentions., only one associated relation definition $d(E_{0},E_{1})$ is given. Here $\mathcal{R}(E_{0},E_{1})$ denotes the whole binary relation space and $d(E_{0},E_{1})$ can be a single sentence or a document specifying the target relation $r(E_{0},E_{1})$. $E_{0}$ and $E_{1}$ are two entity placeholders.
Figure 3: $F_{1}$ (%) scores for different setups on the number and ratio of training instances.
@@
The goal of Definition Only Zero-Shot Relation Extraction task is to extract all relation instances that belong to target relation $r(E_{0},E_{1})$ from any given set of relation instances $\{(s^{i},e_{0}^{i},e_{1}^{i})\}_{i=1}^{n}$ in a binary classification manner. Here $s^{i}$ denotes the $i$-th context text, while $e_{0}^{i}$ and $e_{1}^{i}$ denote two entity mentions inside $s^{i}$. Compared with the traditional zero-shot RE task settings, our task assumes no seen data but only target relation definitions. Our task also assumes the unknown negative relation space while mainstream zero-shot RE assumes known information of all test relations. This makes our task more challenging, but aligns better with real-world scenarios. Further details and discussions are in Appendix A.
2.2 From Few-Shot to Definition-Oriented Low-Resource RE
In this section, we reveal the partial relation pattern coverage issue of commonly adopted RE few-shot setup and further show that the few-shot setup can be converted to our definition-oriented setup while achieving much better results. We take KnowPrompt (Chen et al., 2022b) as the underlying N-way classification model where N equals the number of test relations. It’s a prompt tuning model with robust and strong few-shot performance and does not require hand-crafted prompts.
The LLM derived relation definitions compared with original relation definitions are shown in Appendix G.1 and the experiment results are shown in Fig. 2 and Fig. 7 (in Appendix G.2). The derived relation definitions show that LLM is capable of deducing suitable yet generalizable relation definitions based on few-shot instances. However, the coverage of derived definitions is limited by the coverage of few-shot instances. This is in accordance with our motivation for definition-based low-resource RE setup.
Fig. 2 and Fig. 7 show that our definition derivation and instance generation approach achieves much better performance than the model trained only on few-shot instances. This indicates the approach extends the relation patterns conveyed by the few-shot instances. However, we can see our (15 generated + 4 gold shots) trained model has slightly lower performance than 16 gold shots trained model which is due to the partial coverage of relation semantics conveyed by the 4 gold shots instances. This further illustrates the importance of capturing actual relation definitions instead of few-shot data for low-resource RE approaches.
2.3 Effect of More Initial Seeds
As recent LLMs have enabled larger context windows, a naïve method for improving framework performance is to directly query LLMs for more generated instances. So we conduct such trials on our definition only zero-shot RE task (Sec. 3.1) to investigate whether more initially generated seeds can bring more benefits. The experiments are based on one split of the DefOn-FewRel and leverage GPT-4 as synthesis backbone with 3 random seeds. The positive instances are generated with prompt templates shown in Table 6, Table 7, and Table 8 while the negative instances are randomly sampled from the unlabeled corpus. The instances are used to train a RE model adapted from roberta-large-mnli Liu et al. (2019). The quantitative results are shown in Fig. 5, Fig. 6, and Fig. 3. We can see that synthetic data by LLM is generally beneficial, but generating more initial seeds does not guarantee better results. Larger p&n [^3] We abbreviate the number of positive instances for each target relation as p and the number of negative instances as n. or larger n:p ratio both lead to higher precision and lower recall.
One explanation for these trends is that more positive seeds does not bring larger positive pattern coverage but results in over-fitting. However, more randomly sampled negative seeds lead to larger negative pattern coverage. Therefore the model tends to give positive predictions on instances with some dominant positive patterns but identify other minority patterns as negative. Note that when n&p are small, usually recall is pretty high while precision is low. This may also be caused by our task assumption on unknown negative relation space. The above observations and analysis motivates our design of a feedback-driven follow-up generation and refinement paradigm for instance synthesis.
@@ 3 Methodology
Our proposed REPaL is an iterative framework which consists of three major components: (1) Definition-based seed construction with LLM and the unlabeled corpus (Sec. 3.1). (2) Pattern learning with RE-specialized SLM (Sec. 3.2) which trains a SLM extractor with initial seeds for inference efficiency and performance. (3) Feedback-driven instance improvement and bias rectification (Sec. 3.3) by providing LLMs with independent feedback and leveraging LLMs’ multi-turn conversations for followup positive and negative instance synthesis. The framework overview is shown in Fig. 4.
@@ 3.1 Definition-Based Seed Construction with LLM and Unlabeled Corpus
LLMs, such as the GPT family, are pre-trained for the domain adaptation ability (Radford et al., 2019). They have demonstrated to contain factual relation knowledge (Petroni et al., 2019) and are widely used as data generator for various tasks such as text classification (Meng et al., 2022) and dialogue systems (Abdullin et al., 2024). Follow-up evaluation studies have shown that LLMs are relatively skilled at constrained content generation, storytelling, and rationale generation (Sun et al., 2023; Keskar et al., 2019). Following such observations as well as the goal to tackle the data scarcity issue of zero-shot RE, REPaL first leverages LLMs to generate $K_{p_{0}}$ positive seeds, $\{(s_{p}^{i},e_{0,p}^{i},e_{1,p}^{i})\}_{i=1}^{K_{p_{0}}}$, conditioned on the target relation definition $r(E_{0},E_{1})$. We carefully design three prompts for this step (shown in Appendix B) to encourage the LLM to generate comprehensive patterns from three perspectives in terms of complexity: brief, medium, and implicit. Based on exploratory experiments, they yield different types of relation patterns in accordance with our design expectations. Details of experiments and quantitatively how choice of different prompts affect the results are in Sec. 5.3.
In addition to positive seed generation, constructing negative seeds is also vital for better extraction for the target relation $r(E_{0},E_{1})$, as our task setting assumes no prior knowledge on the negative relation space in order to mimic the real-life deployments (Li et al., 2022). Our negative seed construction is based on the hypothesis that, with a normal large-scale unlabeled corpus, the proportion of target relation instances is relatively small. Thus REPaL conducts random sampling over the unlabeled corpus, which is expected to yield an acceptable set of negative instances denoted as $\{(s_{n}^{i},e_{0,n}^{i},e_{1,n}^{i})\}_{i=1}^{K_{n_{0}}}$. To address potential issues in extreme cases, REPaL designs countermeasures in Sec. 3.3.
@@ 3.2 Pattern Learning with RE-Specialized SLM
Given the relatively high computational costs of fine-tuning and inference with LLMs and the limitation of vanilla in-context learning (ICL) due to LLMs’ difficulties in fully processing long-context demonstrations (Ma et al., 2023), we leverage SLMs as tunable and task-specialized extractors. SLMs also enable faster inference on massive unlabeled corpus to mine feedback in Sec. 3.3. Specifically, we adapt a natural language inference (NLI) model (Obamuyide and Vlachos, 2018; Liu et al., 2019; Sainz et al., 2021) to be relation-specific binary classifiers for simplicity and leave other architectures or LLMs for future research.
@@
where $C,N,E$ denote logits $\mathbf{z}$’s indices for NLI label CONTRADICTION, NEUTRAL, ENTAILMENT respectively. The binary classification loss (Shannon, 1948) for relation $r(E_{0},E_{1})$ is defined as:
@@ 3.3 Feedback-Driven Instance Improvement and Bias Rectification
After obtaining the trained SLM relation extractor, REPaL further introduces a follow-up feedback-driven iterative refinement approach with intuitions listed as follows: (1) The initially generated positive seeds might only have a partial relation pattern coverage or concentrate on common patterns due to LLM’s longtail deficiency (Mallen et al., 2022). Instead of one-time generation, iterative follow-up generations by incorporating feedback from examining previously generated results are better for pattern coverage. (2) Bias or errors can be introduced by previous two steps (Sec. 3.1 and 3.2). One potential source of bias is the generated positive seeds or the randomly sampled negative seeds. Another potential source of bias is the randomness of SLM training over limited training samples (see Fig. 8 and Fig. 9). Identifying and rectifying bias requires a feedback-driven follow-up refinement approach.
Specifically, we first leverage the trained SLM extractor to obtain the inference results on the unlabeled corpus. The first component of the feedback is constructed by randomly sampling instances within a predicted probability range (see Appendix F for details). It is constructed for follow-up positive instance and follow-up negative instance generation respectively, differing only on the probability ranges.
The other component of the feedback is LLMs’ synthesis history. For follow-up positive instance generation, we leverage the multi-turn conversational feature of LLMs where the positive instance synthesis history is prepended to a new turn of conversation asking for positive instances with different patterns. The SLM’s inference feedback is integrated into the new turn of conversation so that LLMs can identify the correct and incorrect patterns learned by SLMs. The double channel feedback design, incorporating both SLM inference and synthesis history, enhances the robustness of this step as the followup synthesis can still rely on LLMs’ self-reflection if the inference feedback is uninformative.
For follow-up negative instance generation, we divide it into two sub-steps following the divide-and-conquer philosophy: the first sub-step is feedback-driven negative relation definition generation and the subsequent sub-step is negative instance generation based on the negative relation definitions. LLM is queried to examine the SLM’s inference feedback and identify the incorrect patterns learned by SLM so that newly generated negative relations can address such patterns. Similarly, for robustness in cases where SLM’s inference feedback is uninformative, we query LLM to generate near-miss negative relation definitions based on the positive relation definition to better distinguish hard cases. After acquiring the negative relation definitions, the second sub-step is the same as the initial positive seed generation process in Sec. 3.1. The prompts used in this section are placed in Appendix C. After obtaining all the follow-up positive and negative instances, we iteratively train the SLM extractor.